Three individuals form a partnership and agree to divide the profits equally. X invests $40,000. Y invests $35,000. Z invests $25,000. If the profits are $2,700, how much less does X receive than if the profits were divided in proportion to the amount invested?
Answer by dragonwalker(73)

Do you mean if the profits are in relation to the amount invested rather than divided equally?
If so then you just want to find the proportion of the total that x invests.
i.e. his investment divided by total investment.
40000/100000 = 0.4
So this is the proportion of the profits that he would receive if divided fairly and the amount is calculated by multiplying this proportion by total profits:
0.4 x 2700 = $1080

However, as profits are divided equally between three people the amount he would get is:

2700/3 = $900

Therefore the difference between the two is 1080 - 900 = $180